# Quantum correlations and metrological advantage among Unruh-DeWitt detectors in de Sitter spacetime

### v1

@@ IV Results and analysis
In this section, we investigate the impact of GH decoherence on the QFI and LQU of scalar fields in dS-ST. To achieve this, we examined freely falling UDW detectors in dS-ST, which interact weakly with a massless scalar field in the dS-invariant BDV. This analysis allowed us to evaluate the degradation of QFI and LQU due to the GH effect, as a function of various physical parameters, including GH temperature $T_{GH}=1/\beta$, the initial state preparation, $\alpha$-vacua, and the detector energy spacing $\omega$.
@@ IV.1 QFI
First, we evaluate how the QFI depends on various parameters to assess its effectiveness as a diagnostic tool for understanding the dS-ST structure. To illustrate our results, we will provide plots of QFI of GH versus the inverse of temperature $\beta$ and different relevant system parameters, particularly the choice of the initial state, $\alpha$-vacua, and the detector energy spacing.
@@
Figure 2 highlights intriguing features of the QFI as a function of the inverse GH temperature, $\beta$, and the initial state parameter, $\tau$, with non-Bunch-Davies (non-BD) vacuum sectors fixed at $|\alpha|=6$. QFI values are presented for fixed detector energy spacings $\omega=3,5,10$. For all cases, a local peak emerges as $\beta$ decreases, suggesting that optimal precision for estimating $\beta=2\pi l$ is achieved with detectors at smaller curvature radii ($l$). QFI reaches a global maximum at an intermediate value $\beta_{int}$, after which it asymptotes to a smaller value. Notably, when $\tau<0$ (e.g., $\tau=-2,-1$), QFI peaks are sharper, increasing and then rapidly decreasing around $\beta_{int}$. The location of $\beta_{int}$ shifts with increasing $\omega$, altering the QFI peak’s position and amplifying the QFI, thereby lowering $\beta_{int}$ in non-BD sectors. This indicates that optimal estimation for $\beta$ depends significantly on the initial state choice. Additionally, as shown in Fig. 2, QFI initially rises to a maximum and then declines as $\beta$ increases, while continuing to increase with $\tau$. This trend underscores that higher precision in estimating $\beta$ is achievable with an appropriate selection of $\tau$.
@@
Figure 3: The plots of QFI among UDW detectors as a function of $\beta$ for various fixed parameters. Panels 3 and 3 depict QFI for different fixed values of $\omega$, namely $\omega=1$ (solid blue), $\omega=3$ (dashed orange), $\omega=5$ (solid red), and $\omega=10$ (dashed green). In panel 3, we set $\tau=-2$ and $|\alpha|=6$, while in panel 3, $\tau=1$ and $|\alpha|=6$. Panels 3 and 3 show QFI for different fixed values of $|\alpha|$, namely $|\alpha|=1$ (solid blue), $|\alpha|=3$ (dashed orange), $|\alpha|=5$ (solid red), and $|\alpha|=10$ (dashed green). In panel 3, we set $\tau=-2$ and $\omega=10$, while in panel 3, $\tau=1$ and $\omega=10$. Figure3 presents the estimation degree of the GH decoherence parameter, $\beta$, as determined by the QFI for various values of $\omega$ and $|\alpha|$. Figure 3 displays that increasing the energy spacing, $\omega$, leads to an enhancement in the estimation precision of $\beta$. Additionally, a larger $\omega$ shifts the maximum bounds of the QFI towards lower values of $\beta$, which means higher precision of parameter estimation can be achieved for the detector with a larger GH temperature. In line with the results presented in Fig. 2, increasing $\tau=1$ results in a significant increase in the maximum QFI. Conversely, the impact of the coherence parameter, $|\alpha|$, on the QFI for $\beta$ is explored in Fig. 3 and Fig. 3 for $\tau=-2$ and $\tau=1$, respectively. The QFI exhibits a single peak that shifts towards lower values of $\beta$ as $|\alpha|$ increases. Notably, the maximum QFI remains relatively constant at approximately $15$ for $\tau=-2$, while it significantly increases to around $70$ for $\tau=1$.
As evidenced in Fig. 2 and 3, a distinct Gaussian trend emerges. Examining the impact of various system parameters on QFI behavior, we observe that an augmentation in the value of $\tau$ results in an elevation of the maximum QFI value. This suggests that a higher $\tau$ enhances the system’s quantum information processing capabilities. Increasing the value of $|\alpha|$ shifts the QFI maximum towards higher GH temperature values, while increasing $\omega$ shifts it towards lower GH temperature values. This indicates that the $\alpha$-vacua and the detector energy spacing influence the probability distribution of quantum system interactions that contribute to enhancing the QFI. Physically, the ascending portion of the Gaussian curve signifies an increase in QFI with rising GH temperature values up to the inflexion point. This implies that a higher GH temperature value promotes quantum interactions that contribute to QFI growth, thereby strengthening quantum coherence among system components and improving quantum information processing capacity. Conversely, the descending portion of the curve, where QFI diminishes with increasing GH temperature, suggests that excessive GH temperature values may disrupt the system’s quantum structure. This disruption can lead to the loss of quantum coherence, consequently diminishing the system’s quantum information processing abilities.
@@ IV.2 LQU
Next, we demonstrate that the LQU arising from $\alpha$-vacua can be harnessed as a resource in RQI tasks. We expect that the amount of LQU will exhibit dependence on the initial state preparation of the detectors and the various selections of superselection sectors of $\alpha$-vacua.
@@
Figure 4 shows the QCs among two accelerated UDW detectors in dS-ST, evaluated using LQU for various initial state parameters $\tau$ and a fixed energy spacing of $\omega=3$. Figure 4 plots the behavior of LQU as a function of $|\alpha|$ with a fixed $\beta=10$. Negative values of $\tau$ yield higher levels of LQU compared to positive values. For negative $\tau$, LQU gradually decreases with increasing $|\alpha|$. However, after reaching $|\alpha|>5.5$, LQU begins to rise again, eventually reaching its maximum limit and evolving to different stable values for different initial states. Conversely, at $\tau=1$, the system’s LQU exhibits a single peak within the range $|\alpha|\in[4,8]$. Outside this range, it remains zero, indicating a transition of the quantum system to a classical state. Similarly, Fig. 4 shows LQU as a function of GH decoherence $\beta$ with $|\alpha|=6$. The system’s behavior resembles that in Fig. 4, although the peak of LQU is shifted to $\beta\in[8,12]$. These figures suggest that transitioning from negative to positive $\tau$ values can diminish LQU, indicating potential decoherence and a shift toward classical behavior. Additionally, the effects of $|\alpha|$-vacua or $\beta$ appear to similarly influence both the generation of LQU and its maximum limits. In particular, the decoherence environment has both positive and negative effects on LQU. Specifically, when the initial state is negative, the LQU increases with the GH effect. This result challenges the widespread belief that decoherence in the environment can only diminish LQU in dS-ST.
@@ V Conclusion
As a result, we have observed the QFI of detectors can achieve the same maximum for different choices of de Sitter-invariant vacuum sectors, given fixed values of $\tau$ and $\omega$. Moreover, the peak maximum of QFI can be enhanced by selecting appropriate energy spacing. However, GH decoherence generally reduces the QFI, which converges to an asymptotic value that is independent of the detector’s initial state preparation. From a metrological perspective, this indicates that enhanced precision in estimating the GH effect can be achieved with relatively high energy spacing. Finally, we emphasize that when the initial state preparation is positive, specifically for $\tau=1$, the related QFI is robust against GH decoherence, in the sense that its local peak can persist over a very long range. This robustness of QFI can significantly facilitate practical relativistic quantum estimation tasks. Similarly, in our LQU analysis, we found that LQU exhibits a striking revival phenomenon as the GH temperature increases. In particular, the amount of LQU can drastically improve after a certain turning point, $\beta_{c}$, for non-BD sectors, due to the competition between GH decoherence and the intrinsic correlation of $\alpha$-vacua. For certain choices of the detectors’ initial state, the measure of LQU will first decrease to zero, then increase, and eventually approach an asymptotic value. Notably, LQU in specific non-BD sectors can persist even under arbitrarily large GH decoherence.

### v2

@@ IV Results and analysis
In this section, we investigate the impact of GH decoherence on the QFI and LQU of scalar fields in dS-ST. To achieve this, we examined freely falling UDW detectors in dS-ST, which interact weakly with a massless scalar field in the dS-invariant BDV. This analysis allowed us to evaluate the degradation of QFI and LQU due to the GH decoherence as a function of various physical parameters, including GH temperature $T_{GH}=1/\beta$, the initial state preparation $\tau$, $\alpha$-vacua, and the detector energy spacing $\omega$.
@@ IV.1 QFI
First, we evaluate how the QFI depends on various parameters to assess its effectiveness as a diagnostic tool for understanding the dS-ST structure. To illustrate our results, we will provide plots of the QFI versus the inverse of temperature $\beta$ and different relevant system parameters.
@@
To gain deeper insights into the influence of the GH temperature and the choice of the initial state $\tau$ on the QFI, we present in Fig. 2 the behavior of the QFI as a function of the parameter $\beta$ and the initial state selection $\tau$, while keeping the non-BDV sectors fixed at $|\alpha|=6$. The QFI values are computed for fixed detector energy spacings, specifically $\omega=3,5,10$. For all examined cases, a local peak in the QFI emerges as $\beta$ decreases, indicating that the optimal precision for estimating $\beta=2\pi l$ is attained when the detectors are positioned at a smaller curvature radius $l$. Furthermore, the QFI exhibits a global maximum at an intermediate value, denoted as $\beta_{\text{int}}$, beyond which it asymptotically decreases. Notably, for $\tau<0$ (e.g., $\tau=-2,-1$), the QFI peaks become more pronounced, rising sharply before undergoing a rapid decline near $\beta_{\text{int}}$. Additionally, the position of $\beta_{\text{int}}$ is observed to shift with increasing detector energy gap $\omega$. This shift modifies the location of the QFI peak while simultaneously enhancing the QFI magnitude, ultimately leading to a decrease in $\beta_{\text{int}}$ within non-BDV sectors. These findings indicate that the optimal estimation of $\beta$ is highly sensitive to the choice of the initial state. Moreover, as depicted in Fig. 2, the QFI initially increases to a maximum before decreasing as $\beta$ grows, while it continues to rise with $\tau$. This trend suggests that higher precision in estimating $\beta$ can be achieved through an appropriate selection of $\tau$. In conclusion, the analysis clearly demonstrates that the QFI is significantly influenced by both the initial state of the UDW detector and the GH temperature.
@@
Figure 3: The plots of QFI among UDW detectors as a function of $\beta$ for various fixed parameters. Panels 3 and 3 depict QFI for different fixed values of $\omega$, namely $\omega=1$ (solid blue), $\omega=3$ (dashed orange), $\omega=5$ (solid red), and $\omega=10$ (dashed green). In panel 3, we set $\tau=-2$ and $|\alpha|=6$, while in panel 3, $\tau=1$ and $|\alpha|=6$. Panels 3 and 3 show QFI for different fixed values of $|\alpha|$, namely $|\alpha|=1$ (solid blue), $|\alpha|=3$ (dashed orange), $|\alpha|=5$ (solid red), and $|\alpha|=10$ (dashed green). In panel 3, we set $\tau=-2$ and $\omega=10$, while in panel 3, $\tau=1$ and $\omega=10$. To elucidate the relationship between the energy spacing $\omega$, different choices of superselection sectors of $\alpha$-vacua, and the QFI of a pair of UDW detectors in dS-ST, we present the QFI of $\eta_{ab}$ as a function of $\omega$ and $\alpha$ in Fig. 3. This analysis is conducted for two distinct cases: (i) when $\tau=-2$, and (ii) when $\tau=1$. Fig. 3 illustrates the estimation precision of the GH decoherence parameter, $\beta$, as quantified by the QFI for various values of $\omega$ and $|\alpha|$. As depicted in Fig. 3, an increase in the energy spacing $\omega$ enhances the estimation accuracy of $\beta$. Furthermore, a larger $\omega$ shifts the peak QFI towards lower values of $\beta$, implying that higher precision in parameter estimation can be achieved for a detector with a greater GH temperature. This result is consistent with the trends observed in Fig. 2, where increasing $\tau=1$ leads to a substantial rise in the maximum QFI. The influence of $\alpha$, which corresponds to different superselection sectors of dS vacua, on the QFI for $\beta$ is further examined in Figs. 3 and 3 for $\tau=-2$ and $\tau=1$, respectively. The QFI exhibits a single peak that shifts towards lower values of $\beta$ as $|\alpha|$ increases. Notably, for $\tau=-2$, the maximum QFI remains approximately constant at $\sim 15$, whereas for $\tau=1$, it undergoes a significant increase, reaching values around $\sim 70$.
As illustrated in Figs. 2 and 3, the behavior of the QFI exhibits a distinct Gaussian trend. A detailed analysis of the dependence of QFI on various system parameters reveals that increasing the value of $\tau$ enhances the maximum QFI. This suggests that QFI is fundamentally influenced by the temperature $\beta$, the energy spacing $\omega$, and different choices of superselection sectors of $\alpha$-vacua. Notably, a larger $\tau$ enhances the system’s quantum information processing capabilities. Furthermore, increasing the parameter $|\alpha|$ shifts the peak of the QFI towards higher values of the GH temperature, whereas an increase in $\omega$ causes the peak to shift towards lower GH temperatures. This implies that both the structure of the $\alpha$-vacua and the detector’s energy spacing modulate the probability distribution of quantum system interactions, thereby contributing to the enhancement of QFI.
@@ IV.2 LQU
Next, we display that the LQU arising from $\alpha$-vacua can be harnessed as a resource in RQI tasks. We expect that the amount of LQU will exhibit dependence on the initial state preparation of the detectors and the various superselection sectors of $\alpha$-vacua.
@@
To gain a deeper understanding of the dependence of QCs between two uniformly accelerated UDW detectors in dS-ST, we analyze the LQU for different initial state choices, denoted by $\tau$, while maintaining a fixed energy spacing of $\omega=3$. Fig. 4 presents the variation of LQU as a function of $|\alpha|$ for a fixed decoherence parameter $\beta=10$. The results indicate that for negative values of $\tau$, the LQU is higher compared to positive $\tau$. In the negative $\tau$ regime, LQU initially decreases with increasing $|\alpha|$. However, beyond $|\alpha|>5.5$, LQU starts to increase again, eventually reaching its maximum limit and stabilizing at different values depending on the initial state. On the other hand, for $\tau=1$, the LQU exhibits a single peak within the interval $|\alpha|\in[4,8]$. Outside this range, LQU vanishes, signifying a transition of the quantum system to a classical state. Similarly, Fig. 4 illustrates the behavior of LQU as a function of the GH decoherence parameter $\beta$, with a fixed value of $|\alpha|=6$. The qualitative behavior remains consistent with that observed in Fig. 4, except that the peak of LQU shifts to the interval $\beta\in[8,12]$. These observations suggest that transitioning from negative to positive values of $\tau$ leads to a reduction in LQU, which implies the onset of decoherence and a shift toward classical behavior. Furthermore, the influence of $|\alpha|$-vacua and the GH decoherence parameter $\beta$ appears to affect both the generation of LQU and its maximum attainable value in a similar manner. Notably, the decoherence environment exhibits both constructive and destructive effects on LQU. Specifically, for negative initial states, LQU increases with the GH effect. This result challenges the conventional belief that environmental decoherence solely leads to a reduction in LQU in dS-ST.
@@ V Conclusion
As a result, we have observed that the QFI of detectors can achieve the same maximum for different choices of de Sitter-invariant vacuum sectors, given fixed values of $\tau$ and $\omega$. Moreover, the peak maximum of QFI can be enhanced by selecting appropriate energy spacing. However, GH decoherence generally reduces the QFI, which converges to an asymptotic value that is independent of the detector’s initial state preparation. From a metrological perspective, this indicates that enhanced precision in estimating the GH effect can be achieved with relatively high energy spacing. Finally, we emphasize that when the initial state preparation is positive, specifically for $\tau=1$, the related QFI is robust against GH decoherence in the sense that its local peak can persist over a very long range. This robustness of QFI can significantly facilitate practical relativistic quantum estimation tasks. Similarly, in our LQU analysis, we found that LQU exhibits a striking revival phenomenon as the GH temperature increases. In particular, the amount of LQU can drastically improve after a certain turning point, $\beta_{c}$, for non-BD sectors due to the competition between GH decoherence and the intrinsic correlation of $\alpha$-vacua. For specific choices of the detectors’ initial state, the measure of LQU will first decrease to zero, then increase, and eventually approach an asymptotic value. Notably, LQU in specific non-BD sectors can persist even under arbitrarily large GH decoherence.
It was pointed out that the qualitative effects of the dS horizon and its associated thermality have been previously reported in the context of black hole spacetimes, particularly in Schwarzschild-de Sitter black holes [80]. Specifically, they suggest that the interplay of multiple horizons (black hole and cosmological) and their associated temperatures can lead to novel entanglement dynamics that deviate from standard asymptotically flat or anti-dS black hole cases. We aim to revisit these issues in future works by investigating the robustness of QFI and LQU under multiple horizons thermality, along with their implications for RQM.